# Outcome-based Reinforcement Learning to Predict the Future

### v3

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has been an effective approach for improving Large Language Models’ reasoning in domains such as coding and mathematics. Here, we apply RLVR methods towards forecasting future real-world events – a challenging task for RL due to the very noisy (and delayed) outcomes involved. Using a novel dataset of recent questions from a prediction market, and accompanying relevant news headlines, we show that a compact (14B) reasoning model can be trained to match or surpass the predictive accuracy of frontier models like o1, while greatly improving probabilistic calibration. The model’s performance is also practically meaningful: in a Polymarket trading simulation, we estimate that its bets would have yielded a return on investment of over 10% across all questions in the test set. We detail and compare approaches used in training our model, including augmenting our training-data with synthetic prediction questions, guardrails for learning stability, and median prediction sampling at inference-time.
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has recently boosted large language model (LLM) performance on benchmarks such as GSM8K, AIME, and MATH(Zhao et al., 2025a; Shao et al., 2024; Guo et al., 2025; Lambert et al., 2024; Xie et al., 2025). RLVR fine-tunes a base model with on-policy reinforcement learning (RL) against objective pass/fail signals, like unit tests or exact numeric solutions(Liu et al., 2023; Su et al., 2025). Complementary self-learning work shows that mixing verifiable non-math corpora into RL pipelines can extend these gains beyond purely symbolic tasks(Akter et al., 2025).
@@
RLVR methods are principally applied to problems whose outcomes are deterministic and instantly verifiable (Murphy, 2025). However, in this work we aim to apply RLVR methods to the problem of forecasting world events. This presents a substantial challenge, due to the inherently noisy and delayed outcomes involved: forecasting requires causal inference, trend extrapolation, and well-calibrated probabilities while supplying only sparse, lagged supervision(Weng, 2024; Qu et al., 2025). Under these conditions, we find that standard GRPO-style updates can drive policies toward extreme overconfidence, gibberish output, or outright training collapse. Adapting RLVR to this setting promises to extend model reasoning ability to an especially demanding domain, potentially unlocking a host of real-world applications.
@@
In this work, we design and empirically validate a stable RL pipeline for probabilistic forecasting, On the algorithmic side, we (i) remove per-question standard-deviation scaling from Group Relative Policy Optimisation (GRPO), (ii) switch to baseline-subtracted advantages in ReMax(Li et al., 2023), and (iii) add lightweight guard-rails, token-length limits, a gibberish filter, and an early-stop criterion, to keep gradients proportional to Brier loss and prevent collapse over more than 100,000 sequential events. On the evaluation side, we collect a novel dataset of 1,265 questions on Polymarket, with accompanying news headlines up to a given prediction date, taking several measures to avoid temporal leakage (a common issue when backtesting the accuracy of forecasting models, see Section 2.1). We assess accuracy with the soft-Brier score and calibration with expected calibration error (ECE). Finally, we quantify economic value by converting each forecast into a set of hypothetical trades and comparing realised profits with those of the frontier reasoning model o1 as a benchmark.
@@
Overall, we find that on the 1,265-question hold-out set, a seven-run ReMax ensemble attains a Brier of 0.190 [0.178, 0.203] and an ECE of 0.062, meaning it achieves the same accuracy as o1 while drastically improving calibration. In a simulated trading scenario, in which the model places a one-share hypothetical bet on every Polymarket question in the test set, ReMax earns a profit of \$52 versus \$39 for o1. Thus, we find that our ReMax ensemble is able to outperform frontier reasoning models in both accuracy and calibration when forecasting future world events, and can leverage this capability to earn meaningful profit in real-money prediction markets.
@@ 2.1 Data and Preprocessing
When developing and evaluating forecasting models, it is critical to avoid sources of ‘temporal leakage’, in which the model is able to use knowledge of future events implicit in its training data or prompt in order to gain an unfair advantage (Paleka et al., 2025). To construct the test set, we follow the same approach as above, but take a variety of additional steps to minimise the possibility of temporal leakage:
@@
    Many questions have indeterminate resolution dates (e.g. “Will X happen some time before Y?’), which could reveal information about the future due to selecting on only resolved questions. To address this, we restrict test-set questions to those which were originally scheduled to close by the time of dataset construction.
@@
Finally, for a second set of experiments, we generate 100,000 additional training questions and answers using Lightning Rod Labs’ proprietary ‘Foresight Learning’ framework. These questions follow a similar format but are created automatically without humans in the loop or labeled data. In the second set of experiments, we use the same initial 10,000 questions with the addition 100,000 synthetic questions mixed in and time-ordered. We reserved the same held-out test set of 1,265 questions for all models.
@@ 2.2 Model
All experiments are conducted with DeepSeek-R1-Distill-Qwen-14B, a 14-billion-parameter model initialised from the open-weight Qwen 2.5-14B base checkpoint (Yang et al., 2024) and further instruction-tuned on the 800k-sample DeepSeek-R1 distilled reasoning corpus(Zhao et al., 2025b). Our fine-tuning regimes take this as the base model and our analyses compare to this base model as well as other benchmarks and frontier models.
@@ 2.3 Reward Design
We treat the forecasting task as a sequential decision problem in which the model receives a textual prompt about a future event, outputs a single probability $\hat{p}\in[0,1]$ that the event will occur, and subsequently observes the binary outcome $y\in\{0,1\}$. The episode reward is the (negative) Brier score(Brier, 1950)
@@
  • Strict Brier: assign the maximum Brier loss of 1.01.01.0 (equivalently a training reward $R=-1$) whenever the regex fails to extract a probability in [0,1][0,1][ 0 , 1 ].
@@
  • Soft Brier: assign a constant Brier loss of 0.250.250.25 (training reward $R=-0.25$) to any malformed or missing forecast, preserving gradient signal while avoiding training collapse.
@@ 2.3.1 RL Algorithms and Updates
First, we test Group Relative Policy Optimization (GRPO)(Shao et al., 2024). For each question $q$ we draw $G$ outputs $\{o^{1},\dots,o^{G}\}$ from the old policy $\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)$ and assign each output $o^{i}$ a scalar reward $r^{i}$. Let
@@
Second, we modify the standard GRPO algorithm by removing the standard-deviation division and set $\hat{A}^{i}=r^{i}-\mu$. This preserves the raw magnitude of especially large forecast errors, potentially improving the model’s ability to correct extreme miscalibrations. Yet, omitting the normalisation can make optimisation more sensitive to outliers, requiring additional guard-rails to mitigate instability. [^1] This vulnerability is partly mitigated in our setting because Brier scores are intrinsically bounded in [0,1][0,1][ 0 , 1 ], which caps the variance of individual rewards and makes the absence of per-question normalisation far less destabilising than it would be for tasks with unbounded numeric losses. This modification parallels the Dr. GRPO method proposed by Liu et al.(Liu et al., 2025).
@@
Third, we also apply ReMax(Li et al., 2023) by first sampling $G$ outputs $\{o^{1},\dots,o^{G}\}$ from the old policy $\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)$, each with reward $r^{i}$. Let $b^{i}$ be a learned baseline for each $o^{i}$. We set the advantage to $\hat{A}^{i}=r^{i}-b^{i}$, removing the need to divide by a standard deviation. The policy is then updated by maximising
@@
Lastly, we also test Direct Preference Optimization(Rafailov et al., 2023), a direct preference-based approach. This method has shown strong performance on QA tasks as well as forecasting tasks (Turtel et al., 2025) and functions as a baseline for some of our analyses.
@@
Across algorithms we keep the optimisation scaffold identical: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$, no weight decay), bfloat16 precision, global grad-norm clip 1.01.01.0, and an entropy bonus coefficient of 0.0010.0010.001. We modify only the levers each method cares about. GRPO uses an actor learning rate of $1\times 10^{-6}$, an initial KL penalty of 0.0050.0050.005, PPO ratio-clip $\epsilon=0.20$, and $G=4$ roll-outs per prompt; Modified-GRPO is identical except that it drops the $\sigma$ division in the advantage to isolate the effect of normalisation. ReMax doubles the actor learning rate to $2\times 10^{-6}$, keeps the KL schedule unchanged, and trains its learned value baseline with $1\times 10^{-6}$ under an MSE loss scaled by 0.50.50.5. DPO is run for 444 epochs at $\beta=0.1$ with a constant $1\times 10^{-5}$ learning rate and a batch size of 128128128 sequences. All runs employ automatic mixed precision and gradient accumulation to emulate two sequences per GPU.
All experiments ran on a single 8-GPU node. The GRPO (10k), ReMax, Modified-GRPO, DPO, and the large-scale GRPO-100k run used eight NVIDIA H100 GPUs.
@@ 2.4 Training Protocol and Stability Measures
Scaling the training to 100k questions introduces stability challenges, particularly related to overconfidence, where some models drift toward 0% or 100% forecasts when per-question reward normalisation is used, and the generation of invalid text or gibberish, where large, noisy datasets can push the policy to produce nonsensical or non-English text. This happens especially when reward gradients fail to distinguish valid forecasts from invalid ones. To maintain stable training at scale, we use a scorer that deducts reward whenever a response violates any of three checks: it flags out-of-context non-English passages, nonsense or random character strings, and the absence of an explanation for the final answer inside a <think>...</think> block. During evaluation the scorer sets the Boolean fields contains_non_english, contains_gibberish, and explains_answer and records the continuous estimates $\mathrm{non\allowbreak\_english\allowbreak\_proportion}$, $\mathrm{gibberish\allowbreak\_proportion}$, and $\mathrm{explanation\allowbreak\_quality}$ (each in [0,1][0,1][ 0 , 1 ]). The downstream reward subtracts penalties such as $\lambda_{\mathrm{lang}}\times\mathrm{non\_english\_proportion}$ and $\lambda_{\mathrm{gib}}\times\mathrm{gibberish\_proportion}$, applies a fixed penalty $\lambda_{\mathrm{miss}}$ when explains_answer is false, and can add up to $\lambda_{\mathrm{exp}}\times\mathrm{explanation\_quality}$ when a rationale is present. The routine also hard-truncates any input beyond 16 000 characters to keep evaluation stable, and its Pydantic schema validation guarantees that any formatting error or missing key zeroes the reward.
@@ 2.5 Failure Modes
We observed a set of different failure modes. For example, in the model trained with GRPO without any guardrails, we find that the model sometimes assigns a probability of 0 to outcomes that are not impossible but rather unlikely. Asigning a probablity of 0 is not something a capable forecaster would do and should be avoided by forecasting models. For example, in response to the question “Will Trump say “Ukraine” 20+ times during Macron presser today?”, the model outputs
@@
Using our Modified GRPO approach, again without any guardrails, we do find that the model is more likely to avoid the extreme overconfidence of the standard GRPO approach, with only a total of 7.9% of predictions landing in either of the extreme buckets. However, we do observe a different failure mode. Some of the answers show language switching between Chinese and English within an output, “Alright, so I need to预测 Patrick Mahomes 在超级碗 LXIX 中是否会 rush 30+码,结果以超过30码为“Over”,否则为“Under”。首先,我会回顾提供的新闻和背景信息,找出相关数据和趋势”, while others show partial segments of ungrammatical phrases and gibberish, such as “Next, I look at the …think that would happen in this interview,” “I s…,” “Lamar Jac…that,” and “They talk about the theme, ticket distribution, and guests, but t…that they’re involved.”
@@ 3.1 Training Algorithm Comparisons
[…]
@@
Soft-Brier and ECE are means across questions; lower is better. Differences are (model – reference). Stars mark two-sided $p$-values for the corresponding pair-wise tests (*<0.05<0.05< 0.05, **∗ ∗<0.01<0.01< 0.01, ***∗ ∗ ∗<0.001<0.001< 0.001).
@@ 3.2 Hypothetical Trading Evaluation
While no model was alone able to achieve a lower Brier score than Polymarket itself, a model’s predictions may nonetheless contain incremental information beyond what is captured by the market price. To provide another test of each model’s value in aiding human prediction, we conduct an simulation analysis in which each model is used to place hypothetical one-share ‘bets’ on all questions in the test set. We then calculate the profit or loss that these bets would have returned on Polymarket.
[…]

### v4

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has been an effective approach for improving Large Language Models’ reasoning in domains such as coding and mathematics. Here, we apply RLVR methods towards forecasting future real-world events – a challenging task for RL due to the very noisy (and delayed) outcomes involved. Using a novel dataset of recent questions from a prediction market, and accompanying relevant news headlines, we show that a compact (14B) reasoning model can be trained to match or surpass the predictive accuracy of frontier models like o1, while greatly improving probabilistic calibration. The model’s performance is also practically meaningful: in a Polymarket trading simulation, we estimate that its bets would have yielded a return on investment of over 10% across all questions in the test set. We detail and compare approaches used in training our model, including augmenting our training-data with synthetic prediction questions, guardrails for learning stability, and median prediction sampling at inference-time. [^1] Dataset available at: https://huggingface.co/datasets/LightningRodLabs/outcome-rl-test-dataset
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has recently boosted large language model (LLM) performance on benchmarks such as GSM8K, AIME, and MATH (zhao2025absolute; shao2024deepseekmath; guo2025deepseek; lambert2024t; xie2025logic). RLVR fine-tunes a base model with on-policy reinforcement learning (RL) against objective pass/fail signals, like unit tests or exact numeric solutions (liu2023rltf; su2025expanding). Complementary self-learning work shows that mixing verifiable non-math corpora into RL pipelines can extend these gains beyond purely symbolic tasks (akter2025nemotron).
@@
RLVR methods are principally applied to problems whose outcomes are deterministic and instantly verifiable (murphy2025rloverview). However, in this work we aim to apply RLVR methods to the problem of forecasting world events. This presents a substantial challenge, due to the inherently noisy and delayed outcomes involved: forecasting requires causal inference, trend extrapolation, and well-calibrated probabilities while supplying only sparse, lagged supervision (weng2024reward; qu2025latent). Under these conditions, we find that standard GRPO-style updates can drive policies toward extreme overconfidence, gibberish output, or outright training collapse. Adapting RLVR to this setting promises to extend model reasoning ability to an especially demanding domain, potentially unlocking a host of real-world applications.
@@
In this work, we design and empirically validate a stable RL pipeline for probabilistic forecasting. On the algorithmic side, we (i) remove per-question standard-deviation scaling from Group Relative Policy Optimisation (GRPO), (ii) switch to baseline-subtracted advantages in ReMax (li2023remax), and (iii) add lightweight guard-rails, token-length limits, a gibberish filter, and an early-stop criterion, to keep gradients proportional to Brier loss and prevent collapse over more than 100,000 sequential events. On the evaluation side, we collect a novel dataset of 1,265 questions on Polymarket, with accompanying news headlines up to a given prediction date, taking several measures to avoid temporal leakage (a common issue when backtesting the accuracy of forecasting models, see Section 2.1). We assess accuracy with the soft-Brier score and calibration with expected calibration error (ECE). Finally, we quantify economic value by converting each forecast into a set of hypothetical trades and comparing realised profits with those of the frontier reasoning model o1 as a benchmark.
@@
Overall, we find that on the 1,265-question hold-out set, a seven-run ReMax ensemble attains a Brier of 0.190 [0.178, 0.203] and an ECE of 0.062, meaning it matches or surpasses o1 in accuracy while markedly improving calibration. In a simulated trading scenario, in which the model places a one-share hypothetical bet on every Polymarket question in the test set, ReMax earns a profit of \$52 versus \$39 for o1. Thus, we find that our ReMax ensemble is able to outperform frontier reasoning models in both accuracy and calibration when forecasting future world events, and can leverage this capability to earn meaningful profit in real-money prediction markets.
@@ 2.1 Data and Preprocessing
When developing and evaluating forecasting models, it is critical to avoid sources of ‘temporal leakage’, in which the model is able to use knowledge of future events implicit in its training data or prompt in order to gain an unfair advantage (paleka2025pitfalls). To construct the test set, we follow the same approach as above, but take a variety of additional steps to minimise the possibility of temporal leakage:
@@
    Many questions have indeterminate resolution dates (e.g. “Will X happen some time before Y?”), which could reveal information about the future due to selecting on only resolved questions. To address this, we restrict test-set questions to those which were originally scheduled to close by the time of dataset construction.
@@
Finally, for a second set of experiments, we generate 100,000 additional training questions and answers using Lightning Rod Labs’ proprietary ‘Foresight Learning’ framework. These questions follow a similar format but are created automatically without humans in the loop or labeled data. Foresight Learning creates training instances from streams of data (e.g. news articles) by using a large language model to generate questions that are difficult to predict at one point in time but easier to verify at a later point in time. In the second set of experiments, we use the same 10,000 questions as the initial experiment plus an additional 100,000 generated questions mixed in and time-ordered. We reserved the same held-out test set of 1,265 questions for all models.
Examples of questions included in this test dataset are “Will Apple launch an iPhone SE on February 19?”, “Will the highest temperature in London be between 37-38°F on February 16?”, and “Will Southampton win on 2025-02-25?”. The full dataset (including questions, relevant news headlines to aid prediction, model reasoning traces and predictions, and the true final resolution) is available here.
@@ 2.2 Model
All experiments are conducted with DeepSeek-R1-Distill-Qwen-14B, a 14-billion-parameter model initialised from the open-weight Qwen 2.5-14B base checkpoint (yang2024qwen2) and further instruction-tuned on the 800k-sample DeepSeek-R1 distilled reasoning corpus (zhao20251). Our fine-tuning regimes take this as the base model and our analyses compare to this base model as well as other benchmarks and frontier models.
@@ 2.3 Reward Design
We treat the forecasting task as a sequential decision problem in which the model receives a textual prompt about a future event, outputs a single probability $\hat{p}\in[0,1]$ that the event will occur, and subsequently observes the binary outcome $y\in\{0,1\}$. The episode reward is the (negative) Brier score (brier1950verification)
@@
  • Strict Brier: assign the maximum Brier loss of $1.0$ (equivalently a training reward R=1R=-1) whenever the regex fails to extract a probability in $[0,1]$.
@@
  • Soft Brier: assign a constant Brier loss of $0.25$ (training reward R=0.25R=-0.25) to any malformed or missing forecast, preserving gradient signal while avoiding training collapse.
@@ 2.3.1 RL Algorithms and Updates
First, we test Group Relative Policy Optimization (GRPO) (shao2024deepseekmath). For each question $q$ we draw $G$ outputs $\{o^{1},\dots,o^{G}\}$ from the old policy $\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)$ and assign each output $o^{i}$ a scalar reward $r^{i}$. Let
@@
Second, we modify the standard GRPO algorithm by removing the standard-deviation division and set $\hat{A}^{i}=r^{i}-\mu$. This preserves the raw magnitude of especially large forecast errors, potentially improving the model’s ability to correct extreme miscalibrations. Yet, omitting the normalisation can make optimisation more sensitive to outliers, requiring additional guard-rails to mitigate instability. [^2] This vulnerability is partly mitigated in our setting because Brier scores are intrinsically bounded in $[0,1]$, which caps the variance of individual rewards and makes the absence of per-question normalisation far less destabilising than it would be for tasks with unbounded numeric losses. This modification parallels the Dr. GRPO method proposed by Liu et al. (liu2025drgrpo).
@@
Third, we also apply ReMax (li2023remax) by first sampling $G$ outputs $\{o^{1},\dots,o^{G}\}$ from the old policy $\pi_{\theta_{\mathrm{old}}}(\cdot\mid q)$, each with reward $r^{i}$. Let $b^{i}$ be a learned baseline for each $o^{i}$. We set the advantage to $\hat{A}^{i}=r^{i}-b^{i}$, removing the need to divide by a standard deviation. The policy is then updated by maximising
@@
Lastly, we also test Direct Preference Optimization (rafailov2023direct), a direct preference-based approach. This method has shown strong performance on QA tasks as well as forecasting tasks (turtel2025llms) and functions as a baseline for some of our analyses.
@@
Across algorithms we keep the optimisation scaffold identical: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$, no weight decay), bfloat16 precision, global grad-norm clip $1.0$, and an entropy bonus coefficient of $0.001$. We modify only the levers each method cares about. GRPO uses an actor learning rate of $1\times 10^{-6}$, an initial KL penalty of $0.005$, PPO ratio-clip $\epsilon=0.20$, and $G=4$ roll-outs per prompt; Modified-GRPO is identical except that it drops the $\sigma$ division in the advantage to isolate the effect of normalisation. ReMax doubles the actor learning rate to $2\times 10^{-6}$, keeps the KL schedule unchanged, and trains its learned value baseline with $1\times 10^{-6}$ under an MSE loss scaled by $0.5$. DPO is run for $4$ epochs at $\beta=0.1$ with a constant $1\times 10^{-5}$ learning rate and a batch size of $128$ sequences. All runs employ automatic mixed precision and gradient accumulation to emulate two sequences per GPU.
All experiments ran on a single 8-GPU node. The GRPO (10k), ReMax, Modified-GRPO, DPO, and the large-scale GRPO-100k run used eight NVIDIA H100 GPUs, for approximately three days. This computational requirement is a substantial constraint for individuals, but is nonetheless a realistic resource available for many computational researchers at academic institutions.
@@ 2.4 Training Protocol and Stability Measures
Scaling the training to 100k questions introduces stability challenges, particularly related to overconfidence, where some models drift toward 0% or 100% forecasts when per-question reward normalisation is used, and the generation of invalid text or gibberish, where large, noisy datasets can push the policy to produce nonsensical or non-English text. This happens especially when reward gradients fail to distinguish valid forecasts from invalid ones. To maintain stable training at scale, we use a scorer that deducts reward whenever a response violates any of three checks: it flags out-of-context non-English passages, nonsense or random character strings, and the absence of an explanation for the final answer inside a <think>...</think> block. During evaluation the scorer sets the Boolean fields contains_non_english, contains_gibberish, and explains_answer and records the continuous estimates $\mathrm{non\allowbreak\_english\allowbreak\_proportion}$, $\mathrm{gibberish\allowbreak\_proportion}$, and $\mathrm{explanation\allowbreak\_quality}$ (each in $[0,1]$). The downstream reward subtracts penalties such as $\lambda_{\mathrm{lang}}\times\mathrm{non\_english\_proportion}$ and $\lambda_{\mathrm{gib}}\times\mathrm{gibberish\_proportion}$, applies a fixed penalty $\lambda_{\mathrm{miss}}$ when explains_answer is false, and can add up to $\lambda_{\mathrm{exp}}\times\mathrm{explanation\_quality}$ when a rationale is present. The routine also hard-truncates any input beyond 16 000 characters to keep evaluation stable, and its Pydantic schema validation guarantees that any formatting error or missing key zeroes the reward.
@@ 2.5 Failure Modes
We observed a set of different failure modes. For example, in the model trained with GRPO without any guardrails, we find that the model sometimes assigns a probability of 0 to outcomes that are not impossible but rather unlikely. Assigning a probability of 0 is not something a capable forecaster would do and should be avoided by forecasting models. For example, in response to the question “Will Trump say “Ukraine” 20+ times during Macron presser today?”, the model outputs
@@
[…]
@@
Figure 1: Mean soft-Brier score (accuracy, left) and mean expected calibration error (ECE, right) for each training algorithm on the Polymarket hold-out set. Error bars show $95\%$ confidence intervals. Lower values are better on both axes.
@@ 3.1 Training Algorithm Comparisons
We measure predictive accuracy with the soft-Brier score, defined as the squared error $(\hat{p}-y)^{2}$ averaged over all 1,265 questions but assigning a soft penalty of $0.25$ whenever a model fails to produce a parseable probability. A score of 0.25 is equivalent to guessing 50% on a question, and thus functions as a suitable stand-in for failed responses. For calibration, we use the expected calibration error (ECE) computed in ten equal-mass probability bins. For both accuracy and calibration, lower scores indicate higher accuracy and better calibration respectively. All statistics are paired across the identical question set; confidence intervals (CI) are two-sided $95\%$ Wald intervals for Brier and bootstrap intervals for ECE unless specified otherwise; every $p$-value reported below is two-sided.
@@
Soft-Brier and ECE are means across questions; lower is better. Differences are (model – reference). Stars mark two-sided $p$-values for the corresponding pair-wise tests (*<0.05<0.05, **<0.01<0.01, ***<0.001<0.001).
@@ 3.2 Hypothetical Trading Evaluation
While no model was alone able to achieve a lower Brier score than Polymarket itself, a model’s predictions may nonetheless contain incremental information beyond what is captured by the market price. To provide another test of each model’s value in aiding human prediction, we conduct a simulation analysis in which each model is used to place hypothetical one-share ‘bets’ on all questions in the test set. We then calculate the profit or loss that these bets would have returned on Polymarket.
For this simulation, we first convert every probability into a one-share trade against the contemporaneous Polymarket price. On Polymarket, users trade shares in forecasting questions, which then pay out a value of \$1 if the question resolves in the direction specified. Therefore, the price of each share corresponds to the market’s estimate of the probability of a given event. For each question in our simulation, we compare the model’s probability $p$ with the last quoted market price $m$: If $p>m$, the simulated betting strategy buys one \$1 long share at $m+0.01$; if $p<m$, it shorts one share at (1m)+0.01(1-m)+0.01; exact ties are broken at random. The added cent approximates fees and slippage. Each trade therefore has a known entry cost $c$, an expected value under the model’s belief ($p$ for longs, 1p1-p for shorts), and a realised value $v\!\in\!\{0,1\}$ at resolution. We evaluate four models: ReMax Ensemble-7, Modified-GRPO, the untuned DeepSeek-R1 base model, and the frontier benchmark OpenAI o1. Profits vcv-c are aggregated in descending order of expected edge to yield cumulative-profit curves, and we report the total return under three bet-selection rules: trading until the edge drops below (i) the model’s own expected calibration error (Edge > ECE), (ii) zero after fees (Edge > 0), and (iii) across all markets.